hartnn (hartnn):

Laplace Transform : f(t) = a sin pt \(0

hartnn (hartnn):

can i do this : \(\Large \int \limits_0^{\pi/p}a \sin (pt) e^{-st} dt +\int \limits _{\pi/p}^{2\pi/p}0 dt\)

hartnn (hartnn):

@mathmale @ganeshie8 @SithsAndGiggles

hartnn (hartnn):

if i do that, i get \(\Large \dfrac{ap}{(p^2+s^2 ) }(1+e^{(-sπ/p)} )\) but i am not sure whether i can split those 2 integrals like i did. If what i did is correct, can someone verify whether my final answer is correct ?

OpenStudy (anonymous):

Ok, what u did there is right so far. Did the question mention anything about that being a periodic pulse? Because then if you take f(x) to be the \[\frac{ 2\pi }{p }\] - periodic extension of fx(s) then your answer would become: \[\frac{ ap(1+e ^{-s \pi/p}) }{ p ^{2}+s ^{2}(1+e ^{-2s \pi/p}) }\]

hartnn (hartnn):

f(t) = f(t+2π/p) implies f(t) is periodic with period 2π/p

hartnn (hartnn):

so my final answer would be \(\Large \dfrac{ ap(1+e ^{-s \pi/p}) }{ (p ^{2}+s ^{2})(1+e ^{-2s \pi/p}) }\) ???

OpenStudy (anonymous):

yep thats right

hartnn (hartnn):

thanks Tushara :)

OpenStudy (anonymous):

no worries

hartnn (hartnn):

oh, isn't it \(\Large \dfrac{ ap(1+e ^{-s \pi/p}) }{ (p ^{2}+s ^{2})(1-e ^{-2s \pi/p}) }\) and the final answer be \(\Large \dfrac{ ap }{ (p ^{2}+s ^{2})(1-e ^{-s \pi/p}) }\) instead ?

OpenStudy (anonymous):

no, the first one is right, from how i remember doing it. Take a look at example 4 on page 4 at http://www.math.vt.edu/people/dlr/m2k_opm_lapper1.pdf that will justify ur answer

hartnn (hartnn):

i have to multiply by \(\Large \dfrac{1}{1-e^{-Ts}}\) right ???

hartnn (hartnn):

for 1st one, there is 1\(\huge +\) ... in the denominator

OpenStudy (anonymous):

I'm curious if writing \(f(t)\) as a Fourier series might be another method...

hartnn (hartnn):

writing it as fourier series and then using s= jw ?

OpenStudy (anonymous):

Sorry I'm not sure what you mean by that

hartnn (hartnn):

hah, nothing. Fourier Transform to Laplace transform, we probably put s=jw....but nvm.. anyways my updated final answer is correct ?

OpenStudy (anonymous):

Oh, I see what you mean. \(j=\sqrt{-1}\), right? I'm used to using \(i\). That might work too, but I'm not sure.

hartnn (hartnn):

yes that. and i have lots of other fourier series examples to solve, so i will not be trying this :P

hartnn (hartnn):

i still think \(\Large \dfrac{ ap }{ (p ^{2}+s ^{2})(1-e ^{-s \pi/p}) }\) should be correct final answer...

OpenStudy (anonymous):

Yeah I'm inclined to agree.

hartnn (hartnn):

thanks! :D